PHOTOIONIZATION OF TMPD IN SOLUTION OBSERVED BY FEMTOSECOND TIME RESOLVED AB-SORPTION SPECTROSCOPY

\underline{JORGE PEON}, TETSURO YUZAWA and BERN KOHLER, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

We have studied the photoionization in solution of the aromatic amine N,N,N’,N’-tetramethyl-p-phenylenediamine (TMPD) in solvents of diverse polarities. Our results characterize the spectral dynamics during the first few picoseconds following monophotonic excitation with 265nm, 100fs pulses using a continuum probe pulse. The 265nm excitation energy (4.7 eV) corresponds to the S_{2} absorption band of TMPD and is below the known gas phase ionization threshold (6.5 eV). As will be discussed, the early evolution of the absorption bands assigned to the S_{1} to S_{n} transition and the radical cation of TMPD provide insight into the role of the solvent in the formation of ions from vibronically unrelaxed species under our experimental conditions.